OpenStudy (anonymous):

The first four terms of a sequence are shown below: 7, 4, 1, -2 Which of the following functions best defines this sequence?

OpenStudy (anonymous):

f(1) = 7, f(n + 1) = f(n) + 3; for n >= 1 f(1) = 7, f(n + 1) = f(n) - 3; for n >= 1 f(1) = 7, f(n + 1) = f(n) - 4; for n >= 1 f(1) = 7, f(n + 1) = f(n) + 4; for n >= 1

OpenStudy (haichi):

C or D

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Would it be D for the +4? because in the sequence it shows it as a positive

OpenStudy (haichi):

its possible

OpenStudy (triciaal):

There are 2 basic sequences the difference is terms is a constant amount -an arithmetic progression is where the constant is an integer added or subtracted and a geometric progression is where the difference in terms is a factor by which we multiply or divide to get the next term for this group we have the first term is 7 and the difference is -3 an AP

OpenStudy (triciaal):

n means which term in the sequence the answer should be B